The length of a swimming pool is 8m longer than its width and the area is 105 meter squared
Answer by Boreal(15235)

width=x
length=x+8 Units in meters; area in sq meters
A=105=x(x+8)
x^2+8x=105
x2+8x-105=0
(x+15)(x-7)=0
x= -15, 7
The width must be positive, so the width is 7m
Length is 7+8=15m
Area is 105 m^2.